OpenStudy (anonymous):

write the equation of a line through the given points in the form Ax + By = C where A B and C are integers and A> 0 (5,4) (-2,6) no idea how to do this. sorry for all the questions. ive got an exam coming up and im having a hard time with the review.

OpenStudy (yuki):

notice that two pairs of (x,y) is given. once you plug them in into the Ax+By =C, you will get a system of equations. Can you solve it from there ?

OpenStudy (anonymous):

no not sure..

OpenStudy (anonymous):

i know the answer is 2x + 7y = 38 but i have no idea how to get there. im really confused.

OpenStudy (yuki):

okay, then do you know that Ax + By = C is called the standard form of a "linear equation"

OpenStudy (yuki):

?

OpenStudy (anonymous):

what do the system of equations look like?

OpenStudy (yuki):

for example, if you plug in (5,4) one of the equations will be 5A+4B = C you do the same with the other set of points

OpenStudy (anonymous):

yes and once i do that im not sure what to do...

OpenStudy (anonymous):

how do i solve that?

OpenStudy (yuki):

you asking me that means that you haven't learned it yet. then I will tell you another way to solve this. to determine a unique line, all you need to two points. since you are given two points, you can find the slope between those points. Once you do that , you are going to plug in one of the points in to this equation y = mx + b called the slope intercept form. x,y is given and m will be the slope. b is the y-intercept and you can solve for it once you plug in all known numbers. after you get y = mx + b, now what you will do is to fix it back into Ax + By = C form by algebraic manipulation.